Instrument Log 

ACSM             C1                  CF Lamont, OK

Date

Technician

Comments

7/7/2016

MG

At 1341 GMT I closed the sampling valve of the ACSM and replaced the Critical orifice. The inlet flow came back close to normal but the incoming RH value was stuck at -6. I then opened and cleaned the cyclone impactor to find it had condensation in it. I then removed the RH sensor on the upstream side of the nafion drier and found corrosion on the sensor. I tried to clean it with contact cleaner but it did nothing for the value. Incoming RH% is still reading -6 and Mentor is sending me a new sensor. RTS 1400 GMT.

7/6/2016

MG

ACSM Critical orifice became clogged and needed cleaned sampling valve was closed at 1330 GMT. ACSM Critical orifice was changed out and sampling valve was re-opened at 1427 GMT.

6/22/16

KT

The instrument computer needed to be rebooted in order for the patch release updates to be installed. At 1410 GMT Instrument mentor rebooted the instrument computer. System is online and operational. RTS at 1425 GMT.

5/31/2016

MG

Pump-box had trouble connecting to to the Turbo pumps through the usb drive. Mentor closed the sample valve and rebooted the computer. RTS 1504

4/18/16

MG

ACSM Critical orifice needed changed. Changed the critical orifice 3 times per mentors request. the o rings that hold it in place will need to be replaced on next change out.

3/1/16

MG

At 22:07 UTC on 2-29-2016 the ACSM Critical orifice became clogged causing the inlet pressure to drop to .8 At 15:06 UTC on 3-1-2016 I swapped out the critical orifice and the inlet pressure returned to the normal of 1.29 I started data acquisition and opened the sample valve. All is well and I will monitor throughout the duration of the day.  RTS 15:06 UTC

2/19/16

MG

Critical Orifice needs to be cleaned. I replaced the Critical Orifice with a new one and am cleaning the old one for a spare. RTS 14:44 GMT

1/20/16

MG

IT department installed software patches on the instrument computer. Instrument mentor tried to reboot computer but it never came back up he then asked me to do a hard reboot. At 1852 GMT I powered down the computer then had to unplug it for several minutes then plugged it back in and powered up the program it came online and began normal operations. RTS at 1852 GMT.

1/4/16

MG

The Central Facility lost power due to an ice storm on Sunday 12-27-2015 around 00:00 GMT. Mentor Restarted computer RTS 1-4-2016 at 15:00 GMT

12/10/15

MG

Mentor was having communications issues with ACSM. Per mentors request, unplugged the USB cables to the aerosol Drier then plugged them back in.

11/18/15

CAM

New servo that was installed last week for the sample/filter valve would not work.  Received instructions from Aerodyne that a replacement IC, which they had shipped with the servo, needed to be installed in the pump control box in order for the new servo to function. Powered the pump control box down and replaced the controller IC.   R/R servo with upgraded replacement model and tested functionality.  Notified mentor that ACSM was ready for operation.  Mentor restarted and RTS at 1836 GMT.

Old servo model HS-645MG.  New servo model HS-7985MG

11/13/15

CAM

Received a new servo from Aerodyne to replace failed unit on sample/filter switching valve.  First connected the new servo to power/control line to test its operation.  New servo would not move in either direction.  Reconnected original servo/valve assembly and although it would not operate at first, after manually moving the valve by hand, it then began to switch properly.  Reconnected the new servo and tested again but it still failed to move.  Left original servo/valve assembly installed and confirmed once more that it would switch properly in both directions.  Notified mentor of situation by voicemail and email.  Did not restart software at this time, awaiting response from mentor.

11/4/15

CAM

The inlet pressure was reading 0.22.  T/S found two different failures. The sample/filter switching valve has failed and cannot be auto-switched.  Manually switched the valve but inlet pressure remained low and unchanged.  Removed the critical orifice and replaced with a new spare.  Inlet pressure then returned to normal, ~1.3 Torr.  Due to failed valve, the instrument will remain down and mentor will order a replacement valve.  Closed sample valve and turned data acquisition off at 1920 GMT. The old critical orifice will be cleaned and returned to spares.

10/6/15

MG

Mentor was here for calibration of ACSM took off-line at 1300 UTC. RTS 1800 UTC.

9/10/15

PD

Notified by Mentor inlet pressure has dropped to 614 after Cyclone impactor was installed.  Mentor asked for Cyclone to be removed for now.  Stopped data acquisition and closed sample valve.  Removed and inspected Cyclone finding no obvious reason for reduced flow.  Blew canned air through Cyclone and reinstalled.  Inlet pressure remained at 620.  Removed Cyclone and connected inlet line.  Started sampling and inlet pressure is now steady at 701, inlet flow steady at 1.3.  RTS at 1445 GMT.

9/9/15

PD

Cyclone impactor was earlier removed at Mentor's request.  After receiving cleaning instructions, impactor was cleaned yesterday, allowed to dry and needs to be installed.  Stopped data acquisition and closed valve.  Installed impactor, opened valve and started data acquisition.  Inlet flow was 1.293 at 1452 GMT during PM rounds and 1.124 after impactor installation.  Will monitor value to see if it increases to normal level of 1.3. 

9/3/15

PD

Mentor suspect Cyclone impactor on ACSM inlet line is partially clogged.  Mentor has a scheduled visit to the site planned for Oct 6th and request impactor removed until then.  Shut sample valve.  Removed Cyclone and connected inlet line with 3/8" Swagelok coupler.  Opened sample valve.  Dryer flow remained unchanged at 2.6 Lpm but inlet flow increased from 1.004 to 1.355 cc/s which is normal operating flow.  Researching how to properly clean impactor before cleaning and reinstalling.  RTS at 1345 GMT.

8/26/15

PD

Measured inlet flow has decreased over the last several days.  Mentor requests the critical orifice replaced.  Shut off inlet sample valve and data collections.  Removed connector to access orifice.  Removed orifice and observed it was bent due to reasons unknown at this time.  Replaced orifice with a used one that had been previously cleaned.  Dryer flow remained the same and inlet flow only slightly increased to 0.980.  Will order a new replacement orifice.  Consulted with Mentor as current and previous instructions on critical orifice installation direction do not agree.  RTS at 2019 GMT.

8/26/15

PD

Computer has lost communication with the pump control box.  Mentor requests powercycling pump control box after he stops program.  After coordinating with Mentor, powercycled pump control box.  Mentor through Radmin was still unable to communicate with pump control box.  Removed USB from from computer that goes to pump box and reconnected.  Connection was then established.  RTS at 1810 GMT.

8/18/15

PD

External vacuum gauge was installed on 8/11/15 at 1754 GMT.  Sample valve was not closed prior to disconnecting vacuum line to install Tee needed for external gauge.  After gauge was installed, flow dropped from a normal of ~1.4 cc/sec down to 0.970 cc/sec.  Mentor suspects critical orifice may have been partially blocked due to removing vacuum line without closing valve first.  Mentor requests new orifice installed.  At 1828 GMT, stopped sample and closed sample valve.  Installed new orifice and opened valve.  Flow rate increased to 0.818 cc/sec.  Removed/reassembled orifice again.  When valve was opened, flow increased to 1.200 then dropped to -0.276, same flow reading observed when valve is closed.  Removed/reassembled orifice again and checked all other connectors in main inlet line.  No connectors were observed as loose and found nothing suspect.  Opened sample valve and flow increased to ~1.4 cc/sec.  Notified Mentor.  RTS at 1938 GMT.

8/11/15

PD

Installing inline vacuum gauges inside AOS building for both the pump directly connected to the ACSM.  At 1754 GMT, cut vacuum lines between instrument and pumps and installed a Tee.  Connected line to gauges.  RTS at 1824 GMT.

6/4/15

KT

Due to a scheduled power shut down the instrument mentor was contacted to advise him of the shut down and requested any action that should be taken. Instrument mentor powered down the ACSM instrument at 1300 GMT. At 1745 GMT instrument powered system on. System is online and operational.

4/27/15

KT

The inlet line had no suction. At 1400 GMT found the inlet vacuum pump not working. Removed and replaced the vacuum pump and suction was restored in the inlet line. RTS at 1415 GMT.

4/22/15

CAM

Received replacement 3-way filter valve and need to install.  Installed valve at 2000 GMT and tested ok for functionality.  Notified mentor instrument was ready for startup.

4/14/15

CAM

Instrument mentor notified Site Ops that the instrument appeared to be stuck in filter mode.  Checked instrument and found the 3-way filter valve had failed.  Instrument was shut down at 1800 GMT and mentor will order a replacement valve.

2/10/15

KT

Instrument mentor requested that the critical orifice be removed and soaked in alcohol overnight then reinstalled into the instrument. At 1510 GMT after the cleaning of the critical orifice was completed it was reinstalled into the instrument. Contacted instrument mentor to let him know that the instrument was online and ready to be started. RTS at 1525 GMT.

1/7/15

KT

The computer appeared to be locked up. Contacted Instrument mentor to have computer issue inspected. Instrument mentor rebooted the computer. RTS at 1434 GMT.

10/24/14

TMG

IT department found corrupt boot loader, which was preventing operating system from loading properly and causing all kinds of problems.  Corrected problem and notified mentor that instrument was ready to be restarted.

10/10/14

CAM

No data has been collected from ACSM since 0807 GMT.  Displays are up on the computer but not updating.  Notified mentor, who requested the Prisma head be power cycled.  Power cycled Prisma at 1540 GMT.  Mentor will restart the instrument remotely.

10/3/14

CAM

Instrument mentor requested the Prisma head be power cycled.  Power cycled Prisma at 1713 GMT.

10/2/14

KT

The computer was not updating. At 1630 GMT power cycled the computer. At 1635 GMT computer came online and the instrument mentor was contacted. RTS at 1645 GMT.

10/1/14

KT

During rounds the computer wasn't updating and the instrument was off. At 1310 GMT reset the UPS. RTS at 1325 GMT.

9/29/14

CAM

Instrument mentor advised that the inlet pressure was not changing when the sampling valve was opened.  Confirmed the 3-way valve for pressure sensor was in the correct position.  Checked sample valve and confirmed it was opening/closing as it should.  Removed 1/8” stainless tubing at the inlet and pressure reading then changed as it should, indicating possible blockage in the 1/8” tube.  Used compressed air to blow through the tube.  Checked and reseated all ultra torr fittings.  Reinstalled 18” tubing and pressure is now behaving and reading as it should.  RTS at 2100 GMT.

9/29/14

CAM

Instrument mentor requested the Prisma be power cycled.  Power cycled Prisma at 1705 GMT.

9/25/14

CAM

New dryer system has been installed and mentor is ready for it to be powered on.  Instrument is then ready for operation.  Powered dryer system on and mentor started software remotely.  ACSM RTS at 1845 GMT. 

9/24/14

CAM

The new dryer system needed to be installed.  Vendor had designed the setup for the dryer to be installed vertically on the ACSM sample inlet, but it could not be installed this way in our situation because the ACSM is installed under a bench.  Re-plumbed the dryer setup to use a 90 degree bend at the ACSM inlet and mounted the dryer horizontally on the backside of the ACSM rack.  Rerouted the 3/8” copper tubing from the stack to the dryer inlet.  Installed the flow controller adjacent to the ACSM rack and made air and electrical connections.  Left voicemail for mentor at 2145 GMT that dryer system was ready for operation, but did not power on at this time.

9/22/14

CAM

The ACSM has returned from the vendor after being repaired and upgraded with a new dryer system.  Installed *ACSM and **computer at 1500 GMT and powered on.  The new dryer system was not installed at this time as per mentor’s request.  Mentor will check instrument remotely and then advise further regarding what to do with the dryer system after conferring with vendor.

*New – 140-103  (WD59483)  **New – C411NS6038 (WD79323)

8/5/14

CAM/KT

Instrument mentor requested the instrument be returned to vendor for repair and upgrades.  Removed *ACSM and **computer at 1600 GMT and took to S/R for shipping. 

*Old - 140-103  (WD 59483)   **Old - C411NS6038 (WD 79323)

8/4/14

KT

SDS could not communicate with the instrument. The main UPS had lost power which caused system to power down. Reset breaker to restore power to the main UPS. The ACSM instrument came online. RTS at 1227 GMT.

7/23/14

CAM

New diaphragm backing pump installed earlier today did not resolve issue of vacuum chamber not pressuring down low enough.  Minimum pressure obtained is approximately 5.6 Torr with the sampling valve closed.  Expected pressure is approximately 1 Torr.  Used methanol soaked cotton swab to test for leaks at 1800 GMT but none found.  Powered pump control box off and removed turbo pump #3 and reseated, ensuring a good seal.  No change in pressure.  Isolated backing pump from turbo pumps / vacuum chamber and pressured dropped to 1.17 Torr as it should.  This confirms backing pump is working properly and leak is somewhere in turbo pump / vacuum chamber.  T/S continues.  Instrument offline at this time.

7/23/14

CAM

New Vacuubrand diaphragm backing pump has arrived and needs to be installed to replace the failing pump.  Powered pump control box off at 1300 GMT. R/R *diaphragm pump.   Original pump power cord had 24V solenoid power piggybacked off the connector so the cords were swapped so the original could be used.  New pump was wired for pulse-width modulation (PWM) control.  Rewired pump for 0-10VDC control   Used mounting feet from the original pump.  Restored pump control box power at 1455 GMT and notified mentor instrument was ready for testing/startup.  Old pump will be rebuilt and used as spare.

*Old - 32592210  *New - 38594705

7/14/14

Mentor

Mentor advised the Vacuubrand diaphragm pump seems to have failed.  System is only pumping down to about 6-7 Torr when it should get down to around 1 Torr.  Mentor placed order for new pump.  Instrument will be down pending pump arrival and replacement.

6/17/14

CAM

Turbo pump #3 has failed and needs to be replaced.  Replacement pump was ordered by mentor and has arrived.  R/R *turbo pump #3 and transferred required fitting from old to new pump.  Powered on pump box and notified mentor at 2015 GMT.   Mentor started vacuum pump and will let run overnight before starting new turbo.  Instrument remains down at this time.  *Old - 229590  *New - IT14201053

5/27/14

CAM

No data had been collected from instrument since the computer was replaced on 5/22.  Mentor had been working with the instrument remotely since that time, but lost comms over the weekend on 5/25.  Found the instrument was without power and completely shut down. Main UPS in the AOS trailer was down.  Breaker for UPS supply line was tripped.  Reset breaker at 1325 GMT and power was restored.  Mentor will restart the instrument remotely.

5/22/14

CAM/TMG

Mentor supplied a replacement computer that had been upgraded to Windows 7.   Mentor remotely shut the instrument and computer down at 1855 GMT.  R/R *computer and restored power.  Notified mentor the exchange was complete and he will transfer files as necessary and restart the instrument.  RTS at 1905 GMT.

*Old – A1C3NS3003 (WD 59915)  *New – C411NS6038 (WD 79323)

3/4/14

KT

System shut down due to power outage that occurred on 3/2/2014 at 2120 GMT. CF was closed 3/3/2014 due to snow storm. CF opened on 3/4/2014. Power cycled the main UPS in the AOS trailer. System came online at 1340 GMT.

2/28/14

CAM

No data had been collected from the ACSM since 0407 GMT on 2/27.  Logged in remotely and found the following ACSM_DQ dialog box open “Runaway OPC: variable npts exceeded value in sub ParseRB npts=6400”.  Cleared the message and then checked the START box in ACSM Configuration and Control window. Notified mentor of the findings and mentor did a complete restart at 1700 GMT.

2/27/14

KT

The computer appears to be locked up. Was not able to get the Flow value. At 1630 GMT emailed instrument mentor and notified him of the issue and asked if he could look into it. Troubleshooting still in progress. RTS 1645 GMT.

2/18/14

CAM

Instrument mentor requested help troubleshooting an issue with the instrument.  Since 2/8, the data has been flatlining as if a valve were stuck in the filter position.  Checked filter and sample line valves/servos and both are operating as they should. Confirmed vacuum pump was operational.  Mentor later discovered a filament had failed in the instrument and he was able to switch over to the backup filament at 1915 GMT.  Mentor will contact vendor for repair recommendation. 

10/31/13

CAM

No data had been collected from the ACSM since 2107 GMT on 10/30.  Logged in remotely and found it the same way as it was yesterday morning.  A dialog box was open trying to open and Igor text file and computer was very sluggish and plots not updating.  Notified mentor and left things as is so mentor could troubleshoot.  Mentor later requested the Prisma be power cycled.  Power cycled Prisma at 1940 GMT.  Mentor will restart software.

10/30/13

CAM

No data had been collected from the ACSM since 1907 GMT on 10/29.  Logged in remotely and found a dialog box was open as if it were trying to open an Igor text file.  Canceled the dialog box and plots resumed updating at 1210 GMT.  Data collection was successful at 1307 GMT, but failed again on the next collection.  Remote check found ACSM computer very sluggish and plots not updating.  Notified mentor (Tom Watson), who remotely restarted the Prisma at 1800 GMT.

6/13/13

CAM

No data being collected from the ACSM since 6/6.  Found the m/z sacn progress bar was stopped halfway through a scan and a dialog box was open as if it were trying to open an Igor file of some type.  Canceled the dialog box and m/z scan resumed briefly but then stopped again.   Emailed mentor but received automated reply that he is out of the office indefinitely.  Followed restart procedure that had been provided by previous mentor.  Data collection resumed at 2152 GMT.  Notified alternate contacts that were provided in mentor’s automated response.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Powered UPS off at 1730 GMT.  Power was restored at 1920 GMT.  Powered UPS on and verified instrument and computer powered up normally.  Notified mentor that software processes would need to be restarted.  Mentor restarted and data collection resumed at 0352 GMT on 2/28.

2/20/13

CAM

The UPS that the ACSM, as well as a few other things, is plugged into continues to have problems, often kicking over to inverter for no reason.  Removed everything else from the UPS so that the ACSM is now the only thing plugged into it.  UPS was accidentally powered off during the process, resetting the ACSM at 1530 GMT.  The turbo pumps seemed to restart ok but the program stopped updating.  Notified mentor that the program may need restarted. 

12/18/12

PAD

Found ACSM vacuum pump located in shed outside AOS building had power but was not turning.  R/R vacuum pump, RTS at 2035 GMT.  Disassembled removed pump head and could find no reason for pump failure.  Removed pump had been installed on 12/5/12 after previous pump had failed.  Vacuum line running between ACSM and pump is nylon reinforced and line is collapsing.  Vacuum can be detected but flow is being restricted.  At 2208 GMT, R/R collapsed line with hard plastic 1/2" OD line.  Powered up pump.  RTS at 2223 GMT.   

12/5/12

CAM

Instrument mentor advised the instrument appeared to be sampling room air.  T/S found the carbon vane vacuum pump that pulls air through the instrument had failed.  R/R pump and RTS at 2210 GMT.  Noticed that the vacuum line between the pump house and building is partially collapsed.  Will need to replace the line soon.

11/28/12

CAM

The UPS is failing intermittently and needs to be replaced.  Worked with mentor to shut the instrument and pumps down at 1730 GMT.  Powered everything off after the turbo pumps finished braking.  Replaced the UPS and restored power at 1800 GMT.  Notified mentor instrument was ready to be restarted.

11/27/12

CAM

ACSM had not collected since 1707 GMT.  Found the computer was powered off.  Powered computer on at 1823 GMT.  Checked system log for any useful information as to why or when computer shut down, nothing found.  Noticed the UPS had only 3 of 5 battery charge indicators lit.  Ran the UPS through 2 self-tests and both passed.  Will check to see if replacement UPS or batteries are available to definitely eliminate the possibility that the UPS is the problem. Notified mentor that instrument needs to be restarted.

11/26/12

CAM

ACSM had not collected since 1007 GMT on 11/23.  Found the computer was powered off.  Powered computer on and notified mentor that the instrument was ready to be started.  Mentor phoned later and asked that the Prisma controller be power cycled.  Power cycled Prisma and instrument was successfully restarted at 1709 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  Power was restored at 2115 GMT after electrical upgrades were complete.  Notified mentor that instrument was ready to be restarted remotely.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  Notified mentor that instrument was ready to be restarted remotely.

7/30/12

CAM/PAD

The replacement filter valve/servo assembly and Nafion dryer tube have been received and need to be installed.  Installed new valve/servo assembly and Nafion dryer.  Contacted mentor who started pumps at 1615 GMT.  Measured sample flow at instrument inlet to be 2.85 l/min.  Mentor will allow pumps to run overnight and start sampling tomorrow.

7/30/12

PAD

Carbon vein vacuum pump that pulls air through ACSM instrument has failed.  Instrument was not operational at time of pump failure.  Instrument was started after pump was replaced and other repairs made.  See additional CM report for more details.   R/R vacuum pump at 1500 GMT.  RTS at 1515 GMT.

7/18/12

CAM

The pump control box and servos need to be tested after pump box was repaired with new voltage regulator yesterday.  Powered on pump box and contacted mentor.  Mentor logged in remotely and successfully moved the sample line valve, but was unable to move the filter line valve.  Confirmed 5V (4.95) to servos.  Removed filter line servo from valve assembly and attempted to move servo while its gear was unmeshed from the valve gear.  Servo would not move, confirming servo failure.  Noted that filter valve also takes more effort to move by hand than the sample valve does.  Possibly the valve being more difficult to move caused the servo to fail, which in turn led to failure of the original regulator.  Talked to vendor and vendor will put together a new filter valve assembly with servo.  Submitted a PR to purchase new assembly and will replace as soon as it is received.

7/17/12

CAM/JEM

The LM7806 voltage regulator in the pump box has failed.  After discussing and receiving approval from vendor, the decision was made to replace the LM7806 with an LM7805, which we have available in the lab.  Took pump box to ERL and technician removed the LM7806 and replaced with LM7805 voltage regulator. Also installed a 10uF/50V capacitor across the regulator’s 24V input   Measured output to be 4.95V.  Installed pump box back into system and notified vendor and mentor that replacement was successful and ACSM was ready for testing.  Due to time constraints, testing will continue tomorrow.

7/16/12

CAM

The servos that open/close the sample and filter valves have stopped responding to software commands.  T/S found the servos were not receiving 6V from the pump box.  Opened the pump box and found the LM7806 voltage regulator had failed, with pin 1 (24V input) melted off.  Spoke to vendor, who recommended the box be returned for evaluation/repair.  Mentor remotely powered the pumps and instrument down at 1545 GMT.  Closed pump valve and removed *pump box.  Instrument will be down until pump box is repaired.  *Old – (no SN), WD 59483, Electrical Safety E005955

7/12/12

CAM

Mentor continues to see problems with ACSM sample flow and data.  Measured the flow at the ACSM inlet to be over 14 L/min, whereas 3 L/min is expected.  T/S determined the nafion tube must have a significant leak, allowing the vacuum from the dryer portion to pull unregulated air through the sample line.  Removed the nafion dryer tube and connected sample line directly to ACSM inlet.  Measured flow at the inlet to be 1.5 L/min.  Checked critical orifice and found it was clean and clear.  Possibly the orifice is actually a 1.5 L/min instead of 3 L/min, or the flow meter is inaccurate.  A new nafion dryer tube has been ordered and should be in in about a week.  While troubleshooting the flow, mentor determined she cannot control the sample or filter valves.  Mentor will discuss the issue with vendor and T/S will continue.  RTS at 2230 GMT.

7/12/12

CAM

Mentor requested the ¼” copper tubing between the base of the stack and the ACSM water trap be replaced with 3/8” copper tubing.  Replaced tubing from 1900 to 1930 GMT. 

7/11/12

CAM

Mentor states water is still getting in the sample line and requested a water trap be inserted prior to the nafion dryer tube.  Constructed a water trap using a Swagelok “T” fitting.  The water trap portion on the bottom of the “T” is clear tubing capped on the opposite end.  Installed at 2200 GMT.

7/10/12

CAM

Mentor requested that the nafion dryer tube critical orifice be checked and the HEPA filter replaced.   Checked critical orifice and observed it to be clear.  Cleaned with compressed air to be sure.  R/R HEPA filter with new spare.  RTS at 2200 GMT.  Notified mentor.

7/10/12

CAM

Desiccant that was installed today is already nearly ½ saturated after only a few hours.  Mentor requested the nafion dryer tube be reinstalled.  Removed the desiccant tube from the sample line at 1950 GMT and connected the nafion dryer tube in its place as it was before, without the water trap.  RTS at 1955 GMT.  Notified mentor.

7/10/12

CAM

Desiccant that was installed yesterday was about ¾ saturated this morning.  Removed desiccant from tube at 1305 GMT and replaced with some of our remaining old stock of Envirogel.  Will see how it does overnight.  RTS at 1310 GMT.

7/9/12

CAM

The orange silica gel desiccant that was installed on the inlet line last week has turned clear and needs to be replaced.  Replaced desiccant at 1537 GMT.  This is desiccant that was recently purchased from a new supplier.  If desiccant continues to saturate so quickly, we may need to try something else.

7/5/12

PAD

Instrument mentor requested an inline desiccant tube be placed on the instrument inlet.  Installed desiccant tube filled with orange silica gel desiccant.

6/28/12

CAM

Instrument mentor requested the Nafion dryer be bypassed.  Mentor remotely stopped instrument at 1950 GMT.  Removed Nafion dryer from the sample line.  Capped off the point where the dryer was pulling air in.  Connected sample line after the water trap directly to instrument using conductive tubing.  Notified mentor maintenance was complete and RTS at 2020 GMT.

5/3/12

PAD

AOS building AC unit failed on 5/2/12.  Sampling was stopped on ACSM due to high room temperature.  Temporary repairs were made to AOS AC system this morning.  Mentor was notified and after room cooled down and vaporizer temperature stabilized, mentor remotely started measurements at 1930 GMT.  RTS at 1945 GMT.

5/2/12

PAD

AC unit in AOS building has failed.  Room temperature is 81F at this time.  Consulted with mentor with concerns of instrument running in high room temperatures.  Mentor stopped software and sampling and put pumps in slower speed.  Instrument is still on but not sampling at this time.  RTS at 1505 GMT.

4/3/12

PAD/CAM

Notified by mentor concentration values dropped to zero starting on 3/31/12 at 1900 GMT.  Mentor requests vacuum pump and inlet lines checked for possible failure or blockage.  At 1516 GMT, verified vacuum pump for instrument was operating correctly.  T/S with mentor found flows of 2.5 l/min inlet and 7 l/min on nafion dryer tube.  Both flows are acceptable.  Further T/S found filter valve was in filter position and not moving to allow sample air.  At 1822 GMT after receiving instructions from vendor on proper adjustment of motor to align motor and valve plastic gears, found one of the bolts that attaches motor was not in place and allowed motor to move out of alignment.  Replacement bolt was found and installed.  Motor and valve gears were aligned and valve is now working correctly.  Mentor restarted program after repairs were made.  RTS at 1938 GMT.

3/26/12

PAD

Vacuum pump for instrument had tripped circuit breaker located at AC power box located in pump rack.  Reset breaker and powered up pump.  RTS at 1434 GMT.

2/1/12

PAD

Mentor requests ACSM calibration performed.  Mentor closed sample valve at 1350 GMT.  ACSM was connected to TDMA to perform calibration.  Worked with mentors of both instruments to complete calibration.  After calibration, returned ACSM to normal operating configuration.  RTS at 1640 GMT.

1/5/12

PAD

At 1326 GMT, remotely accessed computer and error was shown for USB to pumps:  USB device not found.  Unable to remotely check pumps.  At 1623 GMT, checked USB connections and found computer was down.  Mentor had remotely accessed and was trying to get USB issue resolved and then was no longer able to access computer.  Powered up computer and called mentor.  Power cycled Prisma at mentor's request.  Mentor started program and computer is now able to access pump control box to monitor pumps.  RTS at 1643 GMT.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.  Power was restored on 8/9/11 at 0211 GMT.  At 1334 GMT, found computer was inactive.  #5 pump had stopped.  Stopped/started all pumps.  Powercycled Prizma controller and powered up computer.  Notified mentor so programs could be restarted.  RTS at 1359 GMT.

8/2/11

PAD

#5 pump showed braking at 1315 GMT.  Shut down and started all pumps.  Pump 5 is now operating normally.  Mentor requests Prisma controller powercycled.  Disconnected power to controller and mentor remotely shut down computer at 1315 GMT.  Connected power to controller and powered up computer.  Mentor started measurements.  RTS at 1345 GMT.

7/26/11

PAD

Mentor suspects a leak in the inlet line that is causing low Mass Load values.  This has been happening since instrument computer was rebooted on 7/18/11.  Pass3 instrument was also installed on 7/18/11.  At 1206 GMT, disconnected inlet line to allow instrument to sample room air.  No significant change in Mass Load values.  Connected line at 1242 GMT.  Flow is good on blower for main air stack, 30 l/min., and vacuum pump for ACSM instrument shows 20 inHg.  At 1500 GMT, removed U-shaped tube at instrument inlet to check o-ring seals.  I accidentally dropped the critical orifice and used the spare to replace it.  Installed critical orifice and installed line.  No change observed in values.  At 1813 GMT, shut down PASS3 instrument and capped inlet line for PASS3.  Reconnected and restarted PASS3 at 1938 GMT.  No change in values.  At 1930 GMT, bypassed nafion dryer on ACSM instrument.  At 2100 GMT, reconnected nafion dryer.  No change in values.  Troubleshooting with mentor at this time.  RTS at 2100 GMT.

7/18/11

PAD

SDS has not collected data since 7/15/11 at 1700 GMT.  Remotely accessed computer at 1349 GMT.  Configuration and Control window would not open to get flow value.  Notified instrument mentor.  At 1800 GMT, mentor had remotely rebooted computer but could not connect to Prisma controller and requested it power cycled.  Found computer was off.  Checked all power supply connections to computer and restarted.  Power cycled Prisma controller and mentor was able to remotely access computer and connect to Prisma.  Mentor restarted program.  RTS at 1815 GMT.

3/23/11

PAD

Inlet flow rate has decreased to 1.060 cc/sec.  Mentor requests inlet line critical orifice removed and cleaned.  Stopped acquisitions and closed inlet valve at 1807 GMT.  Removed and cleaned orifice according to mentor’s instructions.  Installed orifice, opened valve, and started acquisitions at 1923 GMT.  Inlet flow increased to 1.220 cc/sec.

3/15/11

PAD

Software was upgraded yesterday.  Today  at 1250 GMT, error shown:  ACSM DAQ.  Runaway OPC: variable npts exceeded valve in sub Parserb npts = 6400.  Rebooting QM Start via VNC viewer often solves this problem.  Notified instrument mentor. 

3/14/11

PAD

Mentor requests IE calibration performed.  Shut down TDMA system and replaced atomizer solution with NH4NO3. Added 3 vials supplied by mentor to reach particle concentration of 450 pt/ cc.  Mentor took measurements at that concentration, and then concentration was lowered to 287, 100, and 50 pt/ cc.  Measurements were taken at those levels as well.  Installed HEPA filter on inlet line of instrument to get a zero reading.  Reconnected inlet line, computer was rebooted to reestablish comms with pump controller.  System was RTS at 2054 GMT.

12/29/10

PAD

Notified by mentor that the concentrations of all the species dropped to zero on 12/22/10.  Instrument seems to be operating normally, just not showing any concentrations.  Found vacuum pump for ACSM instrument located in pump rack had failed.  R/R pump.  RTS at 2040 GMT.